40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4696     Entry time: Wed May 11 23:02:52 2011
Author: valera 
Type: Update 
Category: ASS 
Subject: Dither angular stabilizitaion system update 

This is what was done in past two days:

- The ETMY and ITMY pitch and yaw dofs are modulated at 40, 44, 42, 46 Hz respectively (oscillator A=30). The c1ass lockin numbers are 12, 14, 27, 29.

- The NAS55I signal is demodulated at the above frequencies. The demodulated I/Q signal phase is set to shift all signal into I-phase. The lockin inputs are bandpassed around respective frequency f with butter("Bandpass",2,f-0.5,f+0.5). The demod signals are then additionally low passed with butter ("Lowpass",4,0.5) so the servo ugf has to be below 0.5 Hz. The servo filter is p:z 0.0001:0.1.

- The ETMY demodulated signal is fed back to ITMY and visa versa.

- With the above 2x2 servo running we moved the input beam PZTs by hand to follow the cavity.

- At the end we offloaded the servo control signals to the SUS biases again by hand.

- The beam spot centering was estimated by unbalancing the ETMY/ITMY pitch/yaw coil combinations intentionally by 5%, which produces 1.3 mm shift of the node, and comparing the response to the residual signals.

- The dof set up currently is: ETMY pitch lockin 12 -> dof2, ITMY pitch lockin 14 -> dof4, ETMY yaw lockin 27 -> dof7, ITMY yaw lockin 29 -> dof9

- The next step is to demodulate the TRY(X) and servo the input beam PZTs

ELOG V3.1.3-